OpenStudy (anonymous):

2lg5-lg(x+2)=1-lg(2x-1)

zepdrix (zepdrix):

\[\Large 2\log5-\log(x+2)\quad=\quad 1-\log(2x-1)\] So we need to remember a couple of important log rules:\[\Large\color{#DD4747 }{\log a-\log b=\log \frac{a}{b}}\]

zepdrix (zepdrix):

\[\Large\color{#3399AA }{b \log (a)=\log(a^b)}\]

zepdrix (zepdrix):

Understand how we can apply the blue rule to the first term, 2log5?

OpenStudy (anonymous):

Is it the change of base law?

zepdrix (zepdrix):

no.. :o it allows us to bring coefficients in front of the log inside as exponents.

OpenStudy (anonymous):

Oh wait,it's the quotient law right?

OpenStudy (kc_kennylau):

@CalvinGordon the base is always 10

zepdrix (zepdrix):

no not the quotient law yet.. can you see the blue one i posted....? :\

OpenStudy (anonymous):

Power law?

zepdrix (zepdrix):

Yes, that's probably what it's called. Do you see how we can apply that to change 2log5?

OpenStudy (anonymous):

Yes

zepdrix (zepdrix):

So it brings the 2 inside as an exponent, right? :o

OpenStudy (anonymous):

So it should be \[\log_{}5 ^{2}\]

zepdrix (zepdrix):

Ok good,\[\Large \log25-\log(x+2)\quad=\quad 1-\log(2x-1)\]

zepdrix (zepdrix):

Looks like we can apply the quotient rule on the left side now, yes?

OpenStudy (anonymous):

Yup

zepdrix (zepdrix):

Something like this? :o\[\Large \log\left(\frac{25}{x+2}\right)\quad=\quad 1-\log(2x-1)\]

zepdrix (zepdrix):

The right side is a little tricky. You need to recall that if your base ever matches the contents of the log, the result is 1. So there is a sneaky way we can rewrite that 1.\[\Large \log10=1\]

zepdrix (zepdrix):

\[\Large \log\left(\frac{25}{x+2}\right)\quad=\quad \log10-\log(2x-1)\]

OpenStudy (kc_kennylau):

@zepdrix you can throw everything except 1 to the left first :p

OpenStudy (anonymous):

Wait,is lg and log the same thing?

zepdrix (zepdrix):

Ya you could do that also :3 maybe simpler that way.

zepdrix (zepdrix):

What is lg...? I assumed it was a typo.. ive never seen lg before.

zepdrix (zepdrix):

Except on cellphones.

OpenStudy (anonymous):

Well,it's all over my textbook,even the question.It's known as the common logarithm.

zepdrix (zepdrix):

Hmm I had to Google it. I guess the `common logarithm` is the same as log with no subscript, base 10. So yah we're ok.

zepdrix (zepdrix):

Unless it's meant to be log base 2. Have you dealt with that at all?

OpenStudy (anonymous):

Nope,could you give me an example?

OpenStudy (kc_kennylau):

@CalvinGordon please give @zepdrix a medal by clicking on the "best answer" button :) \[2\log5-\log(x+2)=1-\log(2x-1)\\\log25-\log(x+2)=\log10-\log(2x-1)\\\log\frac{25}{x+2}=\log\frac{10}{2x-1}\\\frac{25}{x+2}=\frac{10}{2x-1}\\50x-25=10x+20\\40x=45\\x=1.125\]

OpenStudy (kc_kennylau):

This kind of question is just asking you whether you know the log identities and how to use them

OpenStudy (anonymous):

Alright,noted.